# The Role of Inclusion, Control, and Ownership in Workplace AI-Mediated Communication

### 13 Dec 2023 v2

@@ Abstract.
Given large language models’ (LLMs) increasing integration into workplace software, it is important to examine how potential biases they introduce can impact workers. Stylistic biases in the language suggested by LLMs may cause feelings of alienation and result in increased labor for individuals or groups whose style does not match. We examine how such writer-style bias impacts inclusion, control, and ownership over the work when co-writing with LLMs. In an online experiment, participants wrote hypothetical job promotion requests using either hesitant or self-assured auto-complete suggestions from an LLM and reported their subsequent perceptions of inclusion, control, and ownership. We found that the style of the AI model did not impact perceived inclusion. However, individuals with higher perceived inclusion did perceive greater agency and ownership, an effect more strongly impacting participants of minoritized genders. Feelings of inclusion mitigated a loss of control and agency when accepting more AI suggestions.
@@ 1. Introduction
Large language models (LLMs) have the potential to transform workplace communication but can also exacerbate existing societal biases. LLMs are already seeing widespread use in assisting in common workplace tasks like writing emails and authoring documents. Recent scholarship has shown that workers can use LLMs to increase their productivity and produce higher-quality work (Noy and Zhang, 2023). Despite these potential benefits, LLMs also have drawbacks, including the potential to reproduce social biases (Weidinger et al., 2021; Bender et al., 2021). For instance, when LLMs generate text about particular identity groups, they may display social biases in the content and style of the resulting text (Wan et al., 2023), causing representational harms through stereotyping and demeaning language (Weidinger et al., 2021; Shelby et al., 2023). In addition to representational harms, another category of harm and the one we focus on in this work, is quality-of-service harm (Shelby et al., 2023; Bird et al., 2020), in which systems unequally serve different groups, and which may “result in feelings of alienation, increased labor, and service or benefit loss” for people of different identities (Shelby et al., 2023). For example, if an LLM produces text that is stylistically better aligned with one gender group over another, the result is a model that disparately benefits certain users over others.
Such stylistic biases, whether real or perceived by users, can also impact how individuals and groups of end-users feel about and interact with AI tools. Following the previous example, when an LLM’s text style better suits one group over another, this may impact users’ feelings of inclusion when using that system. Inclusion could affect other psychological factors like users’ sense of the human experiences of control and ownership (Ryan and Deci, 2017) during the co-writing process. For example, lower perceptions of inclusion may lead users to accept fewer AI suggestions; prior work has shown the relationship between accepting AI suggestions and feelings of agency and control (Mieczkowski, 2022).
Given the potential importance of feelings of inclusion on people’s use of AI tools, we set out to examine the relationship between writer-style bias and perceptions of inclusion, control, and ownership while co-writing with AI. In particular, we focus here on the potential impact of AI writing assistants that provide auto-complete suggestions as the user is typing text, for example, while writing an email. In this context, we ask the following questions:
@@
More specifically, in this work, we investigate the impact of a specific communication style—assertiveness—on writers and its impact on minoritized gender groups. We build on a rich history examining gender disparities, particularly in the workplace where assertive language can be very consequential (Amanatullah and Morris, 2010; Smith, 2002) and associated with better work outcomes (Mazei et al., 2015; Amanatullah and Morris, 2010). While there are many cultural differences that guide expectations of gender presentation in the workplace, scholars in gender-based communication describe male talk in Western workplaces as being assertive, powerful, and authoritative (Mullany, [n. d.]; Merchant, 2012; Thimm et al., 2008). If writing assistants making suggestions in a more assertive style are not as well-aligned with the communication style of women and members of other minoritized gender groups, the discrepancy in style alignment could cause an additional burden for women and gender minorities by increasing writing effort and task completion time.
This short paper addresses the research questions above using an online experiment where participants completed a hypothetical but common and consequential workplace writing task:writing to a manager to ask for a promotion. Participants received auto-complete suggestions from one of two LLM-powered AI writing assistants that offered suggestions in different styles: assertive or hesitant. We then asked participants about their feelings of inclusion, control, and ownership in the writing experience.
@@ 2. Background
As LLMs become more robust with the ability to inspire ideas, revise text, and generate short stories, the writing assistants powered by large language models are increasingly seen as co-writers. There is a growing body of work studying the interaction between users and these co-writing systems (Lee et al., 2022; Mirowski et al., 2023). Recent scholarship in this area has focused on understanding how writers evaluate and integrate the suggestions provided by LLMs into their cognitive writing processes (Bhat et al., 2023), and how writers proactively engage with system-generated content (Singh et al., 2022).
Along with their potential benefits, however, these models carry potential risks — risks that may not be equally distributed across different user groups. Prior work has categorized harms that include social systems harms, allocative harms, representational harms, quality of service harms, and interpersonal harms (Weidinger et al., 2021; Shelby et al., 2023; Weidinger et al., 2022). In the context of auto-complete suggestions, prior work has shown that societal biases can cause information harms (a subset of social systems harms) by influencing what content is written (Poddar et al., 2023; Jakesch et al., 2023), and even shifting writers’s attitudes, potentially without their awareness (Jakesch et al., 2023; Kidd and Birhane, 2023).
Quality-of-service harms are another concern identified by researchers. These harms reflect developer’s choices and result in unintended performance disparities based on identity (Shelby et al., 2023; Bird et al., 2020). For example, prior work has shown that audio speech recognition (ASR) systems have significantly higher error rates for Black speakers than White speakers (Koenecke et al., 2020). These error rates can have psychological and behavioral impacts on Black users, such as feelings of frustration and disappointment, and can cause users to have to modify their linguistic patterns (Mengesha et al., 2021). Such quality-of-service harms can lead to alienation and result in greater labor burdens and loss of benefits for marginalized communities (Shelby et al., 2023). In this short paper, we extend research on this topic to consider writer-style biases in auto-complete suggestions and their potential to cause differential impacts on inclusion, control, and ownership for different social groups, with a specific focus on gender.
@@ 3. Methods
We addressed our research questions using an online experiment where participants completed a common workplace writing task. We asked participants to write a hypothetical email to their manager asking for a promotion. We chose the promotion task since it is a situation that requires high assertiveness, a context in which minoritized genders face several challenges (Amanatullah and Morris, 2010; Amanatullah and Tinsley, 2013; Mazei et al., 2015). The experiment had three conditions. Participants in two of the treatments received auto-complete suggestions from a language model. In one of these treatments, the model made assertive auto-complete suggestions, and in the other, it made hesitant auto-complete suggestions. A control group of participants wrote without the assistance of any model.
After the writing task, participants answered a demographic questionnaire. Participants in the AI conditions answered an additional post-task questionnaire regarding their perceptions of inclusion, agency, and ownership while writing with the AI assistant. The questions and 5-point Likert scale answers were adapted from previous work (Mieczkowski, 2022; Jakesch et al., 2023). The research design was approved by [anonymized] IRB. The design was preregistered and is available (anonymously) on OSF111https://osf.io/kwbev?view_only=c5639dcc6292482b8bd79f2086b38d24.
@@ 3.1. Writing App
We used a custom experimental platform combining a rich-text editor and a writing assistant, as seen in Figure 1. was made available by researchers (Jakesch et al., 2023). The suggestions work such that when the participant pauses typing, the system displays suggestions that are 20 words or less. The user can then choose to accept the suggestion by pressing tab or right arrow key, or they can ignore the suggestion by continuing to type. Pressing tab accepts the subsequent word in the suggestion, and a user can press multiple times to accept the full suggestion.
To generate the text suggestions for the experiment in real time, we used GPT-4 via an API call with prompts carefully designed to display assertive or hesitant text. To achieve that goal, we inserted the prompt “In a [self-assured/hesitant] manner ask for a raise” before the participants’ written text. After the entered text, we added the affirming instruction “add one sentence continuing the email in a [self-assured/hesitant] manner asking for a raise”. This way, the suggestions maintained the desired style regardless of the text already entered by the participant.
@@ 3.2. Participant Recruitment
We recruited 750 participants from a gender-balanced sample of English-speaking US-based adults on Prolific. The sample size was calculated using a two-way ANOVA with a power of 80% based on a pre-test survey. To ensure response quality, we required participants to have approval ratings greater than 95% on Prolific. We excluded participants who did not state their gender and manually verified that each writing sample was faithful to the writing task. In the end, we had usable data from 738 participants.
Our experiment emphasized two distinct gender categories: individuals with significant societal influence, predominantly men, and minoritized genders who typically encounter societal disparities. While describing our results, we will refer to Women, Non-Binary or third gender, and “preferred to describe” (also known as Self-Described) as WNBSD participants. Of the 738 participants whose data is used, 48.9% identified as men, and 51.1% as WNBSD: 47.7% identified as women, 3% identified as non-binary or third gender, and 0.4% preferred to describe their gender. The majority of the participants were between 25-34 years old (56%), identified as white (61.9%), received a bachelor’s degree (38.2%), reported working a full-time job (59.4%), and reported working entirely in person (40.6%).
@@ 3.3. Measures
AI reliance. Prior work on how users interact with AI writing assistants uses the measure of mutuality to describe the level of interaction the writer has with the AI assistant (Mieczkowski, 2022; Lee et al., 2022). However, in this work, we are interested in the final text output, not the event-based interactions. We introduce a new measure, AI reliance, as we believe this better intuitively captures the construct. AI reliance is the fraction of AI-written characters in the entire text. The measure uses a score from 0 to 1, where 1 means the assistant wrote the entire text (reflecting complete reliance on the assistant) and 0 means the human wrote the entire text (no reliance).
Inclusion. We draw on the concept of self-extension, which explains how objects become integral to one’s self-concept (Belk, 2013), to develop the notion of inclusion in AI-mediated communication. We focus on identity self-extension where the technology is a reflection of the user (Ross and Bayer, 2021; Mieczkowski, 2022). This literature suggests that if a user feels the technology reflects themselves, they will feel included. We therefore measure inclusion with the following statements: the writing assistant was made for people like me, and the writing assistant sounded like I would write myself.
Control. Prior work has shown that feeling in control is not limited to one action or outcome. Feeling in control can encompass engaging in an action, control over the outcome, or both (Singh et al., 2022; Mieczkowski, 2022). For example, some users may feel a lack of control since they cannot choose which suggestions they are seeing (control during the writing process). However, the fact that they can select which suggestions are incorporated into the message can impact their sense of control over the process and the final version (Bhat et al., 2023). We therefore measure by asking participants how much control did you feel over the process of writing the message and how much control did you feel over the final version of the message (Mieczkowski, 2022).
Ownership. One aspect of ownership involves "mineness"– the idea that a specific target belongs to an individual. The target can include physical objects but can also encompass intangible objects such as words and thoughts. In our experiment, we consider two targets– the message itself and the message style. Therefore, we also investigate ownership over the writing content and style. Our questions included to what extent do you feel like the message you wrote is yours and thinking back on the message writing activity, how much did the message sound like you (Mieczkowski, 2022).
@@ 4. Results
We first provide an overview of the effects of using the AI auto-complete suggestions on the writing task. We examined three components of the writing task: the length of text written by participants, their writing time, and their degree of AI reliance. We performed a two-way ANOVA to compare the means between the three treatment groups (self-assured AI writing style, hesitant AI writing style, and control) and two gender groups (men and WNBSD participants). Overall, participants who wrote with the assistant wrote longer emails (M=160.18, SD=82.16) and spent less time on the task (M=303.82, SD=218.53) than participants in the control group (email length M=95.53, SD=35.64; time on task M=363.66, SD=292.68). The differences are statistically significant (length: ANOVA F=67.86, p= ; time: ANOVA F=6.35, p=0.0018). Post-hoc tests did not observe any differences between the two AI conditions in email length and time on task. In contrast, we did observe significant impacts of AI writing style on AI reliance: participants who wrote with the self-assured writing style model relied on the model more (M=0.55, SD=0.34) than participants who wrote with the hesitant writing style model (M=0.45, SD=0.35). These differences were also significant (t=3.16, p=0.0017). We did not observe gender effects or interaction effects between gender and the AI treatment on any of these metrics.
@@ 4.1. Inclusion
We examine the impact of AI writing style on our key measures, beginning with perceptions of inclusion. Figure 2 shows the frequency of different survey responses for each gender condition (different columns) for the inclusion statements: the writing assistant was made for people like me (“Assistant Was Made for You” on the left), and the writing assistant sounded like I would write myself (“Assistant Sounds Like You” on the right). In the condition with the hesitant AI writing style (Figure 2a, left), 65.1% of men (first column) and 63% of WNBSD participants (second column) either agreed or strongly agreed (blue and dark blue bars in the column) that the writing assistant was made for people like them. As the figure also shows, in the self-assured AI writing style group (right half), 68.3% of the men and 68.6% of WNBSD participants either agreed or strongly agreed with this statement. In general, participants more readily agreed with the “made for you” statement (Figure 2a) compared to the “sounds like you” statement (Figure 2b).
In Figure 2b, of the participants who received the hesitant treatment (left half), 42.8% of the men and 45.7% of the WNBSD participants either agreed or strongly agreed that the assistant sounded like how they would write themselves. In the condition with the self-assured AI writing style, on the right half of the figure, 50% of the men and 51.8% of the WNBSD participants either agreed or strongly agreed. A two-way ANOVA analysis revealed no significant effects for gender, AI treatment group, or the interaction between gender and treatment group across both inclusion questions shown in Figure 2.

### 26 Feb 2024 v3

@@ Abstract.
Given large language models’ (LLMs) increasing integration into workplace software, it is important to examine how biases in the models may impact workers. For example, stylistic biases in the language suggested by LLMs may cause feelings of alienation and result in increased labor for individuals or groups whose style does not match. We examine how such writer-style bias impacts inclusion, control, and ownership over the work when co-writing with LLMs. In an online experiment, participants wrote hypothetical job promotion requests using either hesitant or self-assured auto-complete suggestions from an LLM and reported their subsequent perceptions. We found that the style of the AI model did not impact perceived inclusion. However, individuals with higher perceived inclusion did perceive greater agency and ownership, an effect more strongly impacting participants of minoritized genders. Feelings of inclusion mitigated a loss of control and agency when accepting more AI suggestions.
@@ 1. Introduction
Large language models (LLMs) have the potential to transform workplace communication but can also exacerbate existing societal biases. LLMs are already seeing widespread use in assisting in everyday workplace tasks like writing emails and authoring documents. Recent scholarship has shown that workers can use LLMs to increase their productivity and produce higher-quality work (Noy and Zhang, 2023). Despite these potential benefits, LLMs also have drawbacks, including the potential to reproduce social biases (Weidinger et al., 2022; Bender et al., 2021). For instance, when LLMs generate text about particular identity groups, they may display social biases in the content and style of the resulting text (Wan et al., 2023), causing representational harm through stereotyping and demeaning language (Weidinger et al., 2022; Shelby et al., 2023). Beyond representational harms, we focus in this work on another category of harm: quality-of-service harm (Shelby et al., 2023; Bird et al., 2020), in which systems unequally serve different groups, which may “result in feelings of alienation, increased labor, and service or benefit loss” for people of different identities (Shelby et al., 2023). For example, if an LLM produces text that is stylistically better aligned with one gender group over another, the result is a model that disparately benefits certain users over others.
Such stylistic biases, whether real or perceived by users, can impact how individuals and groups of end-users feel about and interact with AI tools. When an LLM’s text style better suits one group over another, such mismatch may impact users’ feelings of inclusion when using that system. Inclusion could affect other psychological factors like users’ sense of the human experiences of control and ownership (Ryan and Deci, 2017) during the co-writing process. For example, lower perceptions of inclusion may lead users to accept fewer AI suggestions; prior work has shown the relationship between accepting AI suggestions and feelings of agency and control (Mieczkowski, 2022). Given the potential importance of feelings of inclusion on people’s use of AI tools, we set out to examine the relationship between writer-style bias and perceptions of inclusion, control, and ownership while co-writing with AI.
In particular, we focus here on the potential impact of AI writing assistants that provide auto-complete suggestions as the user is typing text, for example, while writing an email. In this context, we ask the following questions:
@@
More specifically, in this work, we investigate the impact of a specific communication style—assertiveness—on writers and its impact on minoritized gender groups. We build on a rich history examining gender disparities, particularly in the workplace where assertive language can be very consequential (Amanatullah and Morris, 2010; Merchant, 2012) and associated with better work outcomes (Mazei et al., 2015; Amanatullah and Morris, 2010). While there are many cultural differences that guide expectations of gender presentation in the workplace, scholars in gender-based communication describe male talk in Western workplaces as being assertive, powerful, and authoritative (Mullany, 2007; Merchant, 2012; Thimm et al., 2003). If writing assistants making suggestions in a more assertive style are not as well-aligned with the communication style of women and members of other minoritized gender groups, the discrepancy in style alignment could cause an additional burden for women and gender minorities by increasing writing effort and task completion time.
This paper addresses the research questions above using an online experiment where participants completed a hypothetical but common and consequential workplace writing task: writing to a manager to ask for a promotion. Participants received auto-complete suggestions from one of two LLM-powered AI writing assistants, one assistant offering suggestions in an assertive style and the other in a hesitant style. We then asked participants about their feelings of inclusion, control, and ownership in the writing experience.
@@ 2. Background
As LLMs become more robust with the ability to inspire ideas, revise text, and generate short stories, the writing assistants powered by large language models are increasingly seen as co-writers. There is a growing body of work studying the interaction between users and these co-writing systems (Lee et al., 2022; Mirowski et al., 2023). Recent scholarship in this area has focused on understanding how writers evaluate and integrate the suggestions provided by LLMs into their cognitive writing processes (Bhat et al., 2023), and how writers proactively engage with system-generated content (Singh et al., 2023).
Along with their potential benefits, however, these models carry potential risks—risks that may not be equally distributed across different user groups. Prior work has identified several categories of harms including social systems harms, allocative harms, representational harms, quality of service harms, and interpersonal harms (Shelby et al., 2023; Weidinger et al., 2022). Prior work has pointed to the existence of such harms in the context of auto-complete suggestions. For example, societal biases in the suggestions can cause information harms (a subset of social systems harms) by influencing what content is written (Poddar et al., 2023; Jakesch et al., 2023), and even shifting writers’ attitudes, potentially without their awareness (Jakesch et al., 2023; Kidd and Birhane, 2023). While we do not know of work showing representational harms in auto-complete environments, given the potential for content influence (Jakesch et al., 2023; Kidd and Birhane, 2023) and the known representational biases that exist in LLM text generation (Shelby et al., 2023; Weidinger et al., 2022), the issue is likely to persist in the auto-complete context as well.
Quality-of-service harms are another concern identified by researchers. These harms reflect developer’s choices and result in unintended performance disparities based on identity (Shelby et al., 2023; Bird et al., 2020). For example, prior work has shown that audio speech recognition (ASR) systems have significantly higher error rates for Black speakers than White speakers (Koenecke et al., 2020). These error rates can have psychological and behavioral impacts on Black users such as feelings of frustration and disappointment; they can cause users to have to modify their linguistic patterns (Mengesha et al., 2021). Such quality-of-service harms can lead to alienation and result in greater labor burdens and loss of benefits for marginalized communities (Shelby et al., 2023). In this short paper, we extend research on this topic to consider writer-style biases in auto-complete suggestions and their potential to cause differential impacts on inclusion, control, and ownership for different social groups, with a specific focus on gender.
@@ 3. Methods
We addressed our research questions using an online experiment where participants completed a common workplace writing task. We asked participants to write a hypothetical email to their manager asking for a promotion. We chose the promotion task since it is a situation that requires high assertiveness in context in which minoritized genders face several challenges (Amanatullah and Morris, 2010; Amanatullah and Tinsley, 2013; Mazei et al., 2015). The experiment had three conditions. Participants in two of the treatments received auto-complete suggestions from a language model. In one of these treatments, the model made assertive auto-complete suggestions, and in the other, it made hesitant auto-complete suggestions. A control group of participants wrote without the assistance of any model.
After the writing task, participants answered a demographic questionnaire. Participants in the AI conditions answered an additional post-task questionnaire regarding their perceptions of inclusion, ownership, and agency while writing with the AI assistant. The questions and 5-point Likert scale answers were adapted from previous work (Mieczkowski, 2022; Jakesch et al., 2023). The research design was approved by Cornell’s IRB. The design was preregistered and is available on OSF111https://osf.io/kwbev?view_only=c5639dcc6292482b8bd79f2086b38d24.
@@ 3.1. Writing App
We used a custom experimental platform combining a rich-text editor and a writing assistant, as seen in Figure 1. The platform code was adapted from prior experimental work on auto-complete systems (Jakesch et al., 2023). The suggestions work such that when the participant pauses typing, the system displays suggestions for completion that are 20 words or less. The user can then choose to accept the suggestion by pressing tab or right arrow key, or they can ignore the suggestion by continuing to type. Pressing tab accepts the subsequent word in the suggestion, and a user can press multiple times to accept the full suggestion.
To generate the interactive text suggestions for the experiment, we used GPT-4 via an API call with prompts carefully designed to display assertive or hesitant text. To achieve that goal, we inserted the prompt “In a [self-assured/hesitant] manner ask for a raise” before the participants’ written text. After the entered text, we added the affirming instruction “add one sentence continuing the email in a [self-assured/hesitant] manner asking for a raise”. This way, the suggestions maintained the desired style regardless of the text already entered by the participant.
@@ 3.2. Participant Recruitment
We recruited 750 participants from a gender-balanced sample of English-speaking US-based adults on the crowdsourcing service Prolific. The sample size was calculated using a two-way ANOVA with a power of 80% based on a pre-test survey. To ensure response quality, we required participants to have approval ratings greater than 95% on Prolific. We excluded participants who did not state their gender and manually verified that each writing sample was faithful to the writing task. In the end, we had usable data from 738 participants.
Our experiment considers two distinct gender categories: men, as category that is usually associated with greater societal influence and dominance, and minoritized genders who typically encounter societal disparities. While describing our results, we will refer to Women, Non-Binary or third gender, and “preferred to describe” (also known as Self-Described) as WNBSD participants. Of the 738 participants whose data is used, 48.9% identified as men and 51.1% as WNBSD: 47.7% identified as women, 3% identified as non-binary or third gender, and 0.4% preferred to describe their gender.
The majority of the participants were between 25-34 years old (56%), identified as white (61.9%), received a bachelor’s degree (38.2%), reported working a full-time job (59.4%), and reported working entirely in person (40.6%).
@@ 3.3. Measures
AI reliance. Prior work on how users interact with AI writing assistants uses the measure of mutuality to describe the level of interaction the writer has with the AI assistant (Mieczkowski, 2022; Lee et al., 2022). However, in this work, we are interested in the final text output, not the event-based interactions. We introduce a new measure, AI reliance, as we believe this measure captures the construct more intuitively. AI reliance is the fraction of AI-written characters in the entire text. The measure uses a score from 0 to 1, where 1 means the assistant wrote the entire text (reflecting complete reliance on the assistant) and 0 means the human wrote the entire text (no reliance).
Inclusion. We draw on the concept of self-extension, which explains how objects become integral to one’s self-concept (Belk, 2013), to develop the notion of inclusion in AI-mediated communication. We focus on identity self-extension where the technology is a reflection of the user (Ross and Bayer, 2021; Mieczkowski, 2022). This literature suggests that if a user feels the technology reflects themselves, they will feel included. We therefore measure inclusion with the following statements: the writing assistant was made for people like me, and the writing assistant sounded like I would write myself. Response options ranged from strongly disagree (1) to strongly agree (5).
Control. Prior work has shown that feeling in control is not limited to one action or outcome. Feeling in control can encompass engaging in an action, control over the outcome, or both (Singh et al., 2023; Mieczkowski, 2022). For example, some users may feel a lack of control since they cannot choose which suggestions they are seeing (control during the writing process). However, the fact that they can select which suggestions are incorporated into the message can impact their sense of control over the process and the final version (Bhat et al., 2023). We therefore measure control by asking participants how much control did you feel over the process of writing the message and how much control did you feel over the final version of the message (Mieczkowski, 2022). Response options ranged from not at all (1) to a great deal (5).
Ownership. One aspect of ownership involves "mineness"—the idea that a specific target belongs to an individual. The target can include physical objects but can also encompass intangible objects such as words and thoughts. In our experiment, we consider two targets—the message itself and the message style. Therefore, we also investigate ownership over the writing content and style. Our questions included to what extent do you feel like the message you wrote is yours and thinking back on the message writing activity, how much did the message sound like you (Mieczkowski, 2022). The response options were the same as those for the control questions.
@@ 4. Results
We first provide an overview of the effects of using the AI auto-complete suggestions on the writing task. We examined three components of the writing task: the length of text written by participants, their writing time, and their degree of AI reliance. We performed a two-way ANOVA to compare the means between the three treatment groups (self-assured AI writing style, hesitant AI writing style, and control) and two gender groups (men and WNBSD participants). Overall, participants who wrote with the assistant wrote longer emails (M=160.18, SD=82.16) and spent less time on the task (M=303.82, SD=218.53) than participants in the control group (email length M=95.53, SD=35.64; time on task M=363.66, SD=292.68). The differences are statistically significant (length: ANOVA F=67.86, p<0.001; time: ANOVA F=6.35, p=0.002). Post-hoc tests did not observe any differences between the two AI conditions in email length and time on task. In contrast, we did observe significant impacts of AI writing style on AI reliance: participants who wrote with the self-assured writing style model relied on the model more (M=0.55, SD=0.34) than participants who wrote with the hesitant writing style model (M=0.45, SD=0.35). These differences were also significant (t=3.16, p=0.002). We did not observe gender effects or interaction effects between gender and the AI treatment on any of these metrics.
@@ 4.1. Inclusion
We examine the impact of AI writing style on our key measures, beginning with perceptions of inclusion. Figure 2 shows the frequency of different survey responses for each gender condition (different columns) for the inclusion statements: the writing assistant was made for people like me (“Assistant Was Made for You” on the left), and the writing assistant sounded like I would write myself (“Assistant Sounds Like You” on the right). In the condition with the hesitant AI writing style (left side of Figure 2a), 65.1% of men (first column) and 63% of WNBSD participants (second column) either agreed or strongly agreed (blue and dark blue bars in the column) that the writing assistant was made for people like them. As the figure also shows, in the self-assured AI writing style group (right side of Figure 2a), 68.3% of the men and 68.6% of WNBSD participants either agreed or strongly agreed with this statement. In general, participants more readily agreed with the “made for you” statement (Figure 2a) compared to the “sounds like you” statement (Figure 2b).
In Figure 2b, of the participants who received the hesitant treatment (left), 42.8% of the men and 45.7% of the WNBSD participants either agreed or strongly agreed that the assistant sounded like how they would write themselves. In the condition with the self-assured AI writing style, on the right side of Figure 2b, 50% of the men and 51.8% of the WNBSD participants either agreed or strongly agreed. A two-way ANOVA analysis revealed no significant effects for gender, AI treatment group, or the interaction between gender and treatment group across both inclusion questions shown in Figure 2.